Phase Camera 1: Ring Aligment NI Shift 20 Jan 09

Go down

Phase Camera 1: Ring Aligment NI Shift 20 Jan 09

Post  mdipaolo on Thu Jan 22, 2009 2:43 pm

I report the summary that We did during the Shift.

The goal of this shift is to try the mirror cool down test on the North arm.
We would also like to centre better the CO2 rings.

There are problems debugging North quick unlock so we will adjust North ring centering.

Looking at last position recorded (916440528) we estimate the following TCS mirror movements
-NI UL 551 to -149 (Up) Time: Jan20,09-10:28:55 GPS: 916478950
-NI BR 4225 to 4525 (Left) Time: Jan20,09-10:28:55 GPS: 916478950

There were problems with locking North arm due to wind.
It was necessary to use gipc in order to keep the lock of the cavity.

Made baseline name NI_20Jan2009 Time: Jan20,09-11:22:31 GPS: 916482166

Power of ring set to 11.8 W

*NI ring shutter opened Time: Jan20,09-11:25:25 GPS: 916482340
*NI ring shutter closed Time: Jan20,09-11:31:34 GPS: 916482709
Test aborted because we moved BR the wrong way
-NI BR 4525 to 3925 (Right) Time: Jan20,09-10:28:55 GPS: 916478950

*NI ring shutter opened Time: Jan20,09-11:39:24 GPS: 916483179
*NI ring shutter closed Time: Jan20,09-11:50:01 GPS: 916483816
Ring still seems decentred so we will move the TCS motors


-NI UL from -149 to -1149 (Up) Time: Jan20,09-11:55:40 GPS: 916484155
-NI BR from 3925 to 2925 (Right) Time: Jan20,09-11:56:44 GPS: 916484219

*NI ring shutter opened Time: Jan20,09-12:01:46 GPS: 916484521
spot does not seem to have moved so we will move motors even more up and right


-NI UL from -1149 to -2149 (Up) Time: Jan20,09-12:18:55 GPS: 916485550
-NI BR from 2925 to 1925 (Right) Time: Jan20,09-12:19:54 GPS: 916485609

*NI ring shutter closed Time: Jan20,09-12:23:09 GPS: 916485804
B7p beam became very badly aligned, cavity became unstable so we were forced to switch off.
Cavity is now ulocked
The red spot moved very quickly across the mirror bottom-left to top-right.





cavity relocked Time: Jan20,09-12:30:34 GPS: 916486249

We will move the position back and leave it a while.
-NI UL from -2149 to -1149 (Down) Time: Jan20,09-12:40:11 GPS: 916486826
-NI BR from 1925 to 2925 (Left) Time: Jan20,09-12:41:24 GPS: 916486899

*NI ring shutter opened Time: Jan20,09-12:51:48 GPS: 916487523

The phase image has come back like before but is still not centred.



We will try to move the mirror to intermediate position.
-NI UL from -1149 to -1649 (Up) Time: Jan20,09-13:19:53 GPS: 916489208
-NI BR from 2925 to 2425 (Right) Time: Jan20,09-13:20:31 GPS: 916489246

The red spot found an intermediate position but is not centred.




We will move back 250.
-NI UL from -1649 to -1399 (Down) Time: Jan20,09-13:44:12 GPS: 916490667
-NI BR from 2425 to 2675 (Left) Time: Jan20,09-13:44:32 GPS: 916490687

The ring is now much better centred:




Conclusion.
The ring alignment is very difficult to do. The phase camera image
seems to have a nonlinear response with TCS mirror movement. The
movement of the B7p beam gives cause for concern about the alignment of
the CO2 ring.
However we have managed to centre the negative lens on the mirror. We
are not yet sure if this corresponds to the ring being centred. Next we
must manually centre the central spot on the TCS bench and confirm that
the central spot is well aligned with the CO2 ring.

I report the phase vs Time and the profile of the same phase during the shift:

NI ring shutter opened Time: Jan20,09-11:25:25 GPS: 916482340


NI ring shutter closed Time: Jan20,09-11:31:34 GPS: 916482709

NI ring shutter opened Time: Jan20,09-11:39:24 GPS: 916483179


NI ring shutter closed Time: Jan20,09-11:50:01 GPS: 916483816

NI ring shutter opened Time: Jan20,09-12:01:46 GPS: 916484521


NI ring shutter closed Time: Jan20,09-12:23:09 GPS: 916485804

NI ring shutter opened Time: Jan20,09-12:51:48 GPS: 916487523


NI ring shutter closed Time: Jan20,09-14:36:22 GPS: 916497420

In the last case i can see one peek of phase at 9000 sec.

mdipaolo

Posts : 106
Join date : 2008-11-05

View user profile

Back to top Go down

Back to top

- Similar topics

 
Permissions in this forum:
You cannot reply to topics in this forum